how fast (how many miles per hour) must I run to cover a distance of 3.1 miles in 25 minutes or less?
Answer by Alan3354(69443)

how fast (how many miles per hour) must I run to cover a distance of 3.1 miles in 25 minutes or less?
-----------------
25 mins = 5/12 hours
---
r = d/t = 3.1/(5/12)
r = 7.44 mi/hr